OpenStudy (anonymous):

log2^(5x+3)=log3^(2x+1)

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (anonymous):

i got it to log2^(5x+3)=log3^(2x+1)

OpenStudy (bahrom7893):

log(a)^b = b*log(a)

OpenStudy (anonymous):

so (5x+3)^2=(02x+1)^3 ?

OpenStudy (owlcoffee):

Say we have the logarithm with base "b" of an exponential number: \[\log_{b} a ^{n}\] by definition we can rewrite the "a^n" as an "a" multiplied "n" times. \[\log_{b} (a.a.a.a.a.a.a.a...a)\] there was a property that had this form: \[(\log_{x} Y)+(\log_{x} Z)=\log_{x} (YZ)\] applying that to the logarithm: \[\log_{b} a+\log_{b} a+\log_{b} a+\log_{b} a+\log_{b} a+\log_{b} (a) ... +\log_{b} a\] but they are all the same factor, for example when we have a+a=2a. Then, when we have an addition of the same factor, it's the same as the times we sum the factor and multiply it by the variable. Now, since the Log with base "b" of "a" is added "n" times, then: \[(n)\log_{b} a\] Therefore I just proven the following equality: \[\log_{b} a ^{n} = (n)\log_{b} a\]

OpenStudy (owlcoffee):

applying it to your ptoblem: \[\log_{} 2^{(5x+3)}=\log_{}3^{(2x+1)} \]

OpenStudy (owlcoffee):

and using the formula I proven: \[(5x+3)\log_{}2=(2x+1)\log_{} 3\]

OpenStudy (anonymous):

that looks very weird

OpenStudy (owlcoffee):

Not really, if we divide both members by log2: \[\frac{ (5x+3)\log2 }{ \log2 }=\frac{ (2x+1)\log3 }{ \log2 }\] the "log2" will simplify and if we multiply by the reciprocal "log3" we get: \[\frac{ 1 }{ (2x+1) }(5x+3)=\frac{ (2x+1)\log3 }{ \log2 }(\frac{ 1 }{ (2x+1) })\] And if you apply the fractionary multiplication, you will see that (2x+1) on the right side will simplify and you'll end up with: \[\frac{ 5x+3 }{ 2x+1 }=\frac{ \log3 }{ \log2 }\]

OpenStudy (owlcoffee):

well, it ws not reciprocal "log3" it was reciprocal "2x+1"

OpenStudy (anonymous):

Sorry I had to go somewhere but the answer the book gives is \[\frac{ \log (3/8) }{\log (32/9) }\]

OpenStudy (anonymous):

your book sucks bro -:(

OpenStudy (anonymous):

such a bad form... but, whatever....

OpenStudy (anonymous):

it really does tho

OpenStudy (mathstudent55):

\(\log2^{5x+3}=\log3^{2x+1}\) \((5x + 3)\log2=(2x + 1)\log3\) \(5x\log2 + 3\log2=2x\log3 + \log3\) \(5x\log2 - 2x\log3 = \log3 - 3\log2\) \(x(5\log2 - 2\log3) = \log3 - 3\log2\) \(x = \dfrac{ \log3 - 3\log2}{5\log2 - 2\log3} \)

OpenStudy (anonymous):

that still doesn't give me the answer I'm supposed to get

OpenStudy (mathstudent55):

What answer are you supposed to get?

OpenStudy (anonymous):

log2^(5x+3)=log3^(2x+1) easier solution to this if this is the base of 10 10^(log2^(5x+3))=10^(log3^(2x+1)) 2^(5x+3)=3^(2x+1) solve for x!

OpenStudy (mathstudent55):

Since x is in the exponents on both sides, don't you still need to take log of both sides?

OpenStudy (anonymous):

log(3/8)log(32/9)

OpenStudy (anonymous):

one over the other scroll up to see it if it doesnt make sense

OpenStudy (mathstudent55):

\(x = \dfrac{ \log3 - 3\log2}{5\log2 - 2\log3}\) \(x = \dfrac{ \log3 - \log2^3}{\log2^5 - \log3^2}\) \(x = \dfrac{ \log3 - \log8}{\log32 - \log9}\) \(x = \dfrac{ \log \frac{3}{8}}{\log \frac{32}{9}}\)

OpenStudy (mathstudent55):

I see. I already continued working on my solution until I got in the form you need.

OpenStudy (anonymous):

sounds wierd, but log _(base 32/9) (3/8)

OpenStudy (anonymous):

my bad... this is how you're suppose to do it log2^(5x+3)=log3^(2x+1) log2^(5x+3)-log3^(2x+1)=0 split it 5xlog2-2xlog3 +3log2-log3=0 x(5log2-2log3)= log3+3log2

OpenStudy (mathstudent55):

That's exactly what I did. Look above.

OpenStudy (anonymous):

yup lol